Prove that if π‘Ž|𝑏 and 𝑏≠0, then |π‘Ž| ≀ |𝑏|.


1
Expert's answer
2022-04-14T04:15:52-0400

"a|b\Rightarrow na=b\Rightarrow \left| b \right|=\left| na \right|=\left| n \right|\left| a \right|\\Since\,\,b\ne 0, n\ne 0\Rightarrow n\in \left\{ -1,1,-2,2,... \right\} \Rightarrow \left| n \right|\geqslant 1\\Then\\\left| b \right|\geqslant 1\cdot \left| a \right|=\left| a \right|"